derive integrated rate law for the first order reaction.
1
Expert's answer
2020-04-20T15:08:40-0400

Let the reaction be


"A\overset{k}{\to} product"

Initial amount of of A be "A_\u00b0"

Rate law eqn will be


"\frac{dA}{dt}= -kA"

Seperating variables

We get


"dA= -kAdt"

Integrating it "From \ A_°\ to \ A"

We get


"[lnA\underset{A_°}{\overset{A}{]}}= -kt\\lnA-lnA_°= -kt\\ln\frac{A}{A_°}= -kt\\A= A_°e^{-kt}"

This is the integrated rate law for first order reaction.